A local company produces an erasable programmable read-only memory (EPROM) for several industrial dlients. It has experienced a relatively flat demand of 3,000 units per year for the product. The EPROM is produced at a rate of 12,000 units per year. The accounting department has estimated that it costs $60 to initiate a production run, each unit costs the company $2 to manufacture, and the cost of holding is based on a 0.25 annual interest rate. What is the maximum level of the on-hand inventory of the EPROMs? O734.847 O636.396 519.615 1469.694
Answer #1

Following are the details provided :

Annual demand of EPROM = D = 3000 units

Ratio of daily demand to daily production rate = d/p = 3000/ 12000 = 0.25

Production set up cost = Cs = $60

Annual unit holding cost = Ch = 0.25 x $2 / unit = $0.50

Economic production quantity ( EPQ )

= square root ( 2 x Cs x D / Ch x ( 1 – d/p) )

= square root ( 2 x 60 x 3000/ 0.50 x ( 1 – 0.25 ))

= square root ( 2 x 60 x 3000 / 0.50 x 0.75 )

= 979.79

Maximum level of on hand inventory

= EPQ x ( 1 – d/p )

= 979.79 x ( 1 – d/p)

= 979.79 x 0.75

= 734.842 ( nearest to possible answer 734.847 )

MAXIMUM LEVEL OF ON HAND INVENTORY OF EPROM = 734.847